OpenStudy (anonymous):

solve: 4x-8=9x+2 and 4x-8= -(9x+2)

OpenStudy (mathlegend):

Remember that when solving equations you are trying to get the variable in this case, 'x' by itself.

OpenStudy (mathlegend):

So starting with 4x-8=9x+2 get the 'x' variable to one side of the equal sign

OpenStudy (mathlegend):

Can you do that step?

OpenStudy (anonymous):

4x+4x-8=9x+4x+2 8=13x+2?

OpenStudy (mathlegend):

Okay... well you are making a mistake. Since on the left side the '4x' is positive.... You would subtract that on both sides.

OpenStudy (mathlegend):

-8=5x+2

OpenStudy (anonymous):

okay then you minus 2 from -8?

OpenStudy (mathlegend):

Yes

OpenStudy (anonymous):

so -10=5x

OpenStudy (mathlegend):

Yes

OpenStudy (anonymous):

then divide -10 by 5 which equals -2

OpenStudy (mathlegend):

Correct!

OpenStudy (anonymous):

ahhh okay! but i dont get the other question is it just like this one?

OpenStudy (mathlegend):

Yes, accept they just added in the distribution of the negative sign.

OpenStudy (mathlegend):

4x-8= -(9x+2) So on the right sign just like you distribute a number... you distribute the minus sign.

OpenStudy (mathlegend):

4x-8=-9x-2 Can you solve from there?

OpenStudy (anonymous):

ill try

OpenStudy (mathlegend):

Okay tell me what you get step by step or as a final answer

OpenStudy (anonymous):

8=-5x-2?

OpenStudy (mathlegend):

-8=-5x-2

OpenStudy (mathlegend):

The minus sign on the left must stay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-8+2=-5x

OpenStudy (mathlegend):

However, though if I were you I would always keep my variable: POSITIVE.

OpenStudy (anonymous):

how do you do that!?

OpenStudy (mathlegend):

4x-8= -(9x+2) 4x-8=-9x+2 right? So instead of subtracting 4 from both sides... lets add 9 to both sides.

OpenStudy (mathlegend):

13x-8=-2

OpenStudy (mathlegend):

Then of course you would add 8 to both sides. 13x=6

OpenStudy (mathlegend):

x= \[\frac{ 6 }{ 13 }\]

OpenStudy (mathlegend):

Which is the actual answer... I don't know why the other way gave you a complete different answer.

OpenStudy (anonymous):

hmm thank you! i get it now

OpenStudy (mathlegend):

So to be safe just keep your variables positive. You are welcome! :)